OpenStudy (anonymous):

Solve 5x + 4y = 40 x + 2y = 11

OpenStudy (anonymous):

x=11-2y 5(11-2y)+4y=40 55-10y +4y=40 55-6y=40 -6y=-15 y=2.5

OpenStudy (anonymous):

I guessing that x = 5/2 y = 6

OpenStudy (anonymous):

x=6 and y=5/2